### 15 Dec 2023 v2

# The Impact of AI Innovations on U.S. Occupations

Abstract
AI’s impact has traditionally been assessed in terms of occupations. However, an occupation is comprised of interconnected tasks, and it is these tasks, not occupations themselves, that are affected by AI. To evaluate how tasks may be impacted, previous approaches utilized manual annotations or coarse-grained matching. Leveraging recent advancements in machine learning, we replace coarse-grained matching with more precise deep learning approaches. Introducing the AI Impact (AII) measure, we employ Deep Learning Natural Language Processing to automatically identify AI patents that impact various occupational tasks at scale. Our methodology relies on a comprehensive dataset of 19,498 task descriptions and quantifies AI’s impact through analysis of 12,984 AI patents filed with the United States Patent and Trademark Office (USPTO) between 2015 and 2020. Our observations reveal that the impact of AI on occupations defies simplistic categorizations based on task complexity, challenging the conventional belief that the dichotomy between basic and advanced skills alone explains the effects of AI. Instead, the impact is intricately linked to specific skills, whether basic or advanced, associated with particular tasks. For instance, while basic skills like scanning items may be affected, others like cooking may not. Similarly, certain advanced skills, such as image analysis in radiology, may face impact, while skills involving interpersonal relationships may remain unaffected. Furthermore, the influence of AI extends beyond knowledge-centric regions. Regions in the U.S. that heavily rely on industries susceptible to AI changes, often characterized by economic inequality or a lack of economic diversification, will experience notable AI impact.
1 Introduction
The rapid advancement of Artificial Intelligence (AI) has undeniably created new business opportunities [1] but has also reshaped the labor market [2, 3, 4, 5], simultaneously reducing hiring in non-AI positions and altering the skill requirements of remaining job postings [6]. Exemplifying this phenomenon, the manufacturing sector has witnessed the automation of previously human-intensive assembly line tasks, while chatbots and virtual assistants have taken over routine inquiries and support functions in customer services [7]. Recent AI advances, including generative AI, may further (re)shape occupations over the long term, fuelling growth in certain sectors and eroding others [8]. AI automation has not only streamlined processes but has also generated economic benefits [9], enabling companies to allocate resources more effectively and to redirect human capital towards higher-value, creative, and complex tasks—(re)skilling and upskilling their workforce [10]. However, this transformation has given rise to divergent viewpoints, with some scholars arguing for a future characterized by AI occupation displacement and mass unemployment [11, 12], while others posit that the AI revolution has the potential to enhance both productivity and quality of work [13].
Previous literature on AI impact on occupations has primarily focused on two classes of methodologies. The first measures AI impact on occupations. More specifically, it breaks down occupations in a finite set of abilities (e.g., manual dexterity, persuasion) and measure the impact on those abilities. Two examples that illustrate this approach are the Frey and Osborne’s [14] method and the AI Occupational Exposure (AIOE) method [4]. Frey and Osborne’s method uses 9 abilities extracted from O*NET database, covering 70 occupations that were manually labeled and extended with a classifier to 702, including roles like clergy, dentists, and chief executives. In contrast, AIOE uses 52 abilities derived from the O*NET database, focusing on 10 Electronic Frontier Foundation (EFF) applications such as image and speech recognition, and language modeling. However, both of these methods share a common limitation. Their reliance on coarse-grained abilities in the computation of AI impact may not fully capture the nuances of AI. Consider, for example, the ability of information ordering. Methods based on abilities may categorize tasks that involve organizing information in the same way, without distinguishing between the highly structured and complex information ordering required for database design and the simpler, routine information tasks of librarians such as alphabetizing files [15].
The second class measures AI impact on tasks rather than occupations. This concept is illustrated by Brynjolfsson et al.’s Suitability for Machine Learning (SML) method [3], which measures the impact of AI using a comprehensive set of 18,156 tasks spanning 964 different occupations. However, SML relies on the assessments of crowdworkers to determine the suitability of specific tasks for machine learning. This reliance may introduce subjective biases from annotators (e.g., varying levels of expertise or cultural factors may lead to inconsistencies), and poses challenges in terms of scalability. Also, similar to Frey and Osborne’s and AIOE, SML is limited by the static nature of its one-time manual labeling. As technology advances and new capabilities emerge such as Large Language Models (LLMs), relying solely on a fixed set of abilities or subjective assessments of tasks suitability for automation becomes increasingly inadequate. For example, a copywriter is likely to be impacted by LLMs [16]. However, if one were to examine copywriters at different points in time, such as in 2010 or 2015, the impact of AI on them would not be constant but would drastically change since language models were not as powerful back in 2010 as they are today. To fully capture the impact of a fast moving technology such as AI, therefore, it is crucial for methods to be adaptable to the ever-evolving technological advancements. To gauge the likelihood of automation, it is essential to identify which systems are poised for construction and commercialization. Annual business surveys serve as a source for measuring the adoption of automation. However, they may be subject to biases (e.g., respondents may over-report positive aspects, prioritize certain business operations, potentially neglecting others, interpret survey questions subjectively without standardized criteria) and are infrequently updated [17]. An alternative, more objective source is patents. Patents are a typical source in scholarly work to identify emerging technological innovations [18, 19, 20]. They provide insights into emerging systems and technologies [21], leading Webb to study AI innovations by comparing occupational task descriptions with patent titles [2]. This method, employing a dictionary approach to identify verb-noun pairs associated with both tasks and patent titles, has limitations though. Another method, similar in its approach to Webb’s term matching, employs a normalized term matching approach to determine the similarity between tasks and patents [22], and does so to the specific area of robotics rather than AI. The method most similar to ours was proposed in [23, 24], in which, using word embedding, patents are matched with broad occupation categories from the American Community Survey. However, that level of categorization is not suitable for researching the characteristics of specific jobs. Overall, this second class of approaches has used either term matching or word embedding. The problem is that term matching does not capture the semantic meaning of words (e.g., it does not distinguish between ‘bank’ as in ‘data bank’ or as in a financial institution) [25], and word embedding does not account for word ordering (e.g., ‘data entry and analysis’ and ‘analysis of entry data’ are considered similar based on word embeddings yet are two different tasks). As a result, these methods either miss relevant patents or return spurious task-patent matches, as detailed in Tables C.1 and C.2.
To overcome these limitations, we introduced and validated the AI Impact (AII) measure. AII utilizes 19,498 task descriptions from O*NET and assesses AI’s impact through innovations found in 12,984 AI patents filed with the United States Patent and Trademark Office (USPTO) from 2015 to 2020. For consistency with prior methods, our analysis extends up to 2020 (additional analyses up to 2022 are in §D). Built on Sentence-T5 (ST5)[26], a natural language processing-focused deep learning framework (Figure 6), the method gauges semantic similarity between occupation task descriptions and patent descriptions (explained in §4.1) by embedding not individual words but the entire document (e.g., the entire patent’s abstract), allowing for considering both semantic meaning and word ordering. The AII score is calculated in three steps. Firstly, the method identifies the most similar patent for each task based on maximum cosine similarity. Secondly, it categorizes a task as AI-impacted if its similarity with the most similar patent surpasses a threshold at the 90th percentile, as previous literature suggested [21] and this work further empirically validated (§B). We select, for each task, the closest patent rather than counting the number of closely linked patents. This approach provides a more targeted understanding of the specific innovations or solutions directly relevant to that particular task, avoiding dilution of the analysis with potentially less pertinent or peripheral patents. Also, given that patents can be general, our approach addresses this by using the text not only in titles, as previous approaches did [2], but also in abstracts, which are more likely to contain application domains or references to specific tasks (illustrated in Table B.1). Finally, the AII score for an occupation is computed by dividing the number of tasks impacted by AI patents by the total tasks for that occupation. For insights into AI’s economic and spatial ramifications, the method aggregates AII scores for each occupation at industry sector- or regional-levels (elaborated in §4.2).
2 Results
2.1 Most- and Least-Impacted Occupations
Table 1: 20 most- and least-impacted occupations ranked by the AII (Artificial Intelligence Impact) measure.
Rank Most-impacted Least-impacted
1 Cardiovascular Technologists and Technicians Pile-Driver Operators
[…]
We compared the most-impacted (highest AII scores) occupations with the least-impacted (lowest AII scores) occupations (Table 1 only reports the 20 most- and least-impacted occupations for brevity and comparability with previous methods), and did so by thematically analyzing the AI patents associated with the tasks of each group’s occupations (as described in §4.5).
Most-impacted Occupations. The highest-impact occupations mainly consist of white-collar occupations such as cardiovascular technicians, orthodontists, health information technicians, numerical and process control programmers, and multi-media artists. Indeed, the AII score, binned by education levels from high school to Master of Sciences, shows that the highest impact is seen for jobs requiring Bachelor’s degrees (Figure 1). Highest-impacted occupations are primarily found in the healthcare, information technology, and manufacturing. Their tasks can be completed in a very specific sequence, and the inputs and outputs of these tasks can be expressed in a machine-readable format. To see how, we examined the types of tasks that patents have automated, and organized the patents into three themes: healthcare, information technology, and manufacturing (Table 2). For the theme of healthcare, from 2015 to 2020, 60% of the tasks done by cardiovascular technologists and 48% of those done by magnetic resonance imaging (MRI) technologists have been impacted by patents automating health records’ management and analyzing MRI scans. In addition to the advanced healthcare occupations, we observed a significant number of patents impacting less skilled healthcare personnel including, for example, patents recording and evaluating patient questionnaires. For the theme of information technology, over the same five years, 47% of tasks done by software developers and 40% of those done by computer programmers have been impacted by patents automating programming tasks and developing workflows. For the theme of manufacturing, over the same five years, 45% of tasks done by truck and tractor operators and 40% of earth drillers’ tasks have been automated. These automated tasks are planning processes such as water-well drilling rigs and driving through electric-powered trucks. More specifically, qualitative analyses based on ethnographic work on long-haul truck driving (perceived as a prime target for AI displacement) have shown that truckers are not being displaced by intelligent systems—the trucker is though still in the cab driving [27]. Instead, they are experiencing the emergence of intelligent systems that monitor their body directly. Such emergence has not been instantaneous but has evolved over the years with the introduction of positioning systems, then smartwatches, and now even more advanced health monitoring devices [28, 29]. The problem is that previous work has assumed that, over those five years, the impact was constant [4]. That was mainly because they could not factor in innovations as they were coming in. By contrast, our score is adjusted based on incoming innovations, and, as such, changes over time.
Figure 1: AII score binned by the level of education: high school, associate degrees from community colleges, bachelor’s degrees, and master’s degrees in science. This binned score was obtained by averaging the scores across all occupations in a given education category, weighted by the total employment for those binned occupations.
Table 2: The most-impacted occupations based on AII scores. Each occupation’s score is calculated as the number of impacted tasks divided by the total number of tasks for that occupation. For each occupation, a task-patent pair is presented, corresponding to the most impacted task by the patent (i.e., the patent with the highest similarity score to the task) which is determined by calculating the textual similarity between the task’s description and a patent’s title plus abstract.
Rank Occupation Sector Task Patent Similarity # Impacted Tasks Total Tasks AII
1
Cardiovascular Technologists and Technicians
Healthcare
Observe gauges, recorder, and video screens […] during imaging of cardiovascular system.
Automated analysis of vasculature in coronary angiograms
0.8753 15 25 0.60
[…]
Orthodontists
Healthcare
Study diagnostic records, such as medical or dental histories […] to develop patient treatment plans.
Patient-Specific Therapy Planning Support Using Patient Matching
0.8460 6 10 0.60
[…]
Medical Records and Health Information Technicians
Healthcare
Plan, develop, maintain, or operate a variety of health record indexes […]
Dynamic health record problem list
0.8350 10 20 0.50
[…]
Numerical Tool and Process Control Programmers
Manufacturing
Determine the sequence of machine operations, […]
Methods and apparatuses for cutter path planning and for workpiece machining
0.8376 8 16 0.50
[…]
Multi-Media Artists and Animators
Information
Create pen-and-paper images to be […] animated by computer.
Line drawing generation
0.8396 7 14 0.50
[…]
Magnetic Resonance Imaging Technologists
Healthcare
Operate optical systems to capture dynamic magnetic resonance imaging (MRI) images […]
MRI system and method using neural network for detection of patient motion
0.8615 11 23 0.48
[…]
Nuclear Medicine Technologists
Healthcare
Gather information on patients’ illnesses and medical history to guide the choice of diagnostic procedures for therapy.
Patient-Specific Therapy Planning Support Using Patient Matching
0.8630 8 17 0.47
[…]
Software Developers, Applications
Information
Design […] mathematical models to predict and measure outcome and consequences of design.
Adaptive predictive analytics for design modification requests
0.8295 7 15 0.47
[…]
Electro-Mechanical Technicians
Manufacturing
Train robots, using artificial intelligence software or interactive training techniques, […]
Architectures, systems and methods for program defined entertainment state system, […]
0.8535 16 35 0.46
[…]
Industrial Truck and Tractor Operators
Manufacturing
Move controls to drive gasoline- or electric-powered trucks, cars, or tractors and transport materials between loading, processing, and storage areas.
Autonomous Truck Unloading for Mining and Construction Applications
0.8301 5 11 0.45
[…]
Sound Engineering Technicians
Arts and entertainment
Record speech, music, and other sounds on recording media, using recording equipment.
Media capture and process system
0.8293 6 14 0.43
[…]
Computer Programmers
Information
Prepare detailed workflow charts and diagrams […]
System and method for generating flowchart from a text document using natural language processing
0.8443 7 17 0.41
[…]
Life, Physical, and Social Science Technicians, All Other
Scientific and technical services
Adjust remotely sensed images for optimum presentation […]
Smart imaging using artificial intelligence
0.8374 27 67 0.40
[…]
Earth Drillers, Except Oil and Gas
Construction
Operate water-well drilling rigs and other equipment […]
Systems and methods for oilfield drilling operations using computer vision
0.8280 12 30 0.40
[…]
Medical Transcriptionists
Healthcare
Transcribe dictation for a variety of medical reports […]
Methods for improving natural language processing […] for automated generation of a clinical summarization report […]
0.8367 6 15 0.40
[…]
Grinding, Lapping, Polishing, and Buffing Machine Tool Setters, Operators, and Tenders, Metal and Plastic
Manufacturing
Move machine controls to index workpieces, and to adjust machines for pre-selected operational settings.
Motor control apparatus
0.8218 7 18 0.39
[…]
Airline Pilots, Copilots, and Flight Engineers
Transportation
Steer aircraft along planned routes, using autopilot and flight management computers.
Cumulative model for scheduling and resource allocation for airline operations
0.8367 10 26 0.38
[…]
Physical Scientists, All Other
Scientific and technical services
Develop automated routines to correct for the presence of image distorting artifacts, such as ground vegetation.
Detection and Management of Target Vegetation Using Machine Vision
0.8345 9 24 0.38
[…]
Computer-Controlled Machine Tool Operators, Metal and Plastic
Manufacturing
Implement changes to machine programs, and enter new specifications, using computers.
Registering collaborative configuration changes of a network element in a blockchain ledger
0.8370 10 27 0.37
[…]
Biomedical Engineers
Healthcare
Diagnose and interpret bioelectric data, using signal processing techniques.
Electrocardiogram Device and Methods
0.8622 7 19 0.37
Figure 2: The number of newly impacted tasks each year for the most affected occupations, combined with the most frequently occurring words in the patents influencing those tasks, are organized around the themes of healthcare, information technology, and manufacturing. These were derived qualitatively and describe the main themes emerging from the patents. In 2016, patents mentioning “patient,” “medical,” and “image” matched tasks in healthcare. In 2017, patents mentioning “robot,” “method,” and “module” matched tasks in manufacturing. In 2018, patents mentioned “machine,” “path,” and “user,” impacting tasks in IT. In 2019, patents mentioning “data,” “learning,” and “model” matched tasks in both IT and Manufacturing.
By analyzing the annual increase in impacted tasks and examining the most frequently occurring words in patents’ abstracts each year (Figure 2), we identified two categories of highly impacted occupations. The first category includes occupations that experienced a sudden impact, while the second category includes occupations that faced continuous impact over time.
Occupations that underwent sudden impacts are predominantly within the healthcare sector. This impact became most pronounced in 2016, when eight new tasks were affected, gradually decreasing to just one new impacted task by 2019. In 2016, patents began to significantly impact healthcare by automating medical imaging and diagnosis through machine learning models, devising treatment plans and medical devices, and recording and analyzing patient data. This impact continued into 2017 and 2018, although to a lesser degree, focusing on predicting optimal radiation therapy doses, dental treatment plans, and processing medical patient data.
Occupations that sustained continuous impact over time are primarily in information technology and manufacturing. In the information technology sector, occupations such as software developers, and in manufacturing, occupations such as earth drillers, saw consistent increases from zero new impacted tasks in 2016 to four new impacted tasks in 2019. In information technology, the impact of patents became noticeable in 2017 when they began training robots to design and execute iterative tests on chemical samples, working on aerial and satellite imagery to create products such as land cover maps, and implementing speech recognition and natural language processing on audio. This impact steadily rose and extended into 2019, with patents integrating machine learning into software systems, automating tasks such as troubleshooting networks and code reviews. In manufacturing, patent impact emerged in 2018, focusing on optimizing supply chain logistics and planning material dumping operations. This impact persisted into 2019, further supporting predictive maintenance and operational optimization such as determining aircraft conditions.
Least-impacted Occupations. The least-impacted occupations mainly consist of blue-collar occupations such as pile-driver operators, agricultural graders and sorters, floor sanders, aircraft cargo handling supervisors, and dredge operators. Indeed, the AII score, binned by education levels, shows that the lowest impact is seen for jobs requiring a high school diploma or less (Figure 1). These occupations are primarily found in agriculture, transportation, accommodation and food services, and construction sectors, where the core tasks and responsibilities revolve around physical and manual labor and typically do not require a wide range of complex mental or physical activities, nor do they involve abstract reasoning. In agriculture, least-impacted tasks involve food inspection and dairy management for livestock. In transportation, least-impacted tasks involve scheduling and resource allocation for airline operations, engine sound control, and vehicle dispatching. In accommodation and food services, tasks include monitoring and recording food temperatures. In construction, least-impacted tasks involve the maintenance and operation of equipment and machinery. In addition to these sectors, by examining occupations with nearly zero AII scores, we found another least impacted set of occupations. Contrary to the previous least impacted occupations which involve manual labor or dexterity, managerial occupations typically require human interactions, having tasks ranging from contract negotiation to proposal review to internal assessments and audits.
Figure 3: Sector-level AII scores for: (a) all sectors; (b, left) sectors with lowest rate of change from 2015 to 2020; and (b, right) sectors with highest rate of change.
2.2 Most- and Least-Impacted Industry Sectors
Having explored the disparities in the impact on occupations, our focus now shifts to understanding the impact on industry sectors. Just as in our previous analysis, where we compared occupations based on AII scores, we have similarly compared industry sectors to identify the most- and least-impacted ones (Figure 3a), and to track AI impact over the years (Figure 3b) (as described in §4.3). By again thematically analyzing the occupations’ descriptions within each sector, we aimed to uncover potential explanations for the most- and least-impacted sectors.
Highly-impacted sectors include healthcare, information technology, and manufacturing, which aligns with our thematic analysis of tasks within the most-impacted occupations. These sectors have experienced a significantly high rate of impact. From the thematic analysis, two possible explanations emerge.
The first explanation lies in the nature of the tasks within these sectors, which are likely to be replaced by AI-enhanced hardware. For example, in healthcare, tasks involving the use of X-rays or MRI scans, such as those used by radiologists, have been automated by patents on advanced medical equipment and devices. Similarly, in manufacturing, tasks that involve the examination of chemical or biological samples, such as those performed by food science technicians, can now be executed by AI-enhanced hardware. The second explanation is that the occupations within these sectors entail tasks demanding extensive data analysis and processing. In information technology, film editors, for example, engage in video data editing, which our patent analysis found to have been streamlined by AI-based software. Likewise, scientists in healthcare, information technology, or manufacturing often handle large volumes of data, and recent patents deal with both structured and unstructured data (e.g., using deep-learning for computational biology [30]).
Least-impacted sectors, on the other hand, include construction, accommodation and food services, real estate, education, public administration, and finance and insurance. These sectors have experienced a significantly low rate of impact. From the thematic analysis, three possible explanations emerge.
The first explanation centers on the fact that occupations within certain sectors are often associated with low-skilled or physical/manual labor. For instance, in construction, tasks range from assembling solar panels to maintaining pipe systems to operating various drills, all of which require physical labor. Additionally, manual dexterity is challenging to automate. In public administration, numerous occupations, such as police officers and firefighters, still require manual skills.
The second explanation is that occupations within certain sectors demand basic and non-specialized skills. For instance, in accommodation services, the tasks of waiters and baristas typically involve minimal specialized knowledge or vocational training, such as serving food and drinks. Similarly, in real estate, brokers or clerks primarily require training in overseeing transactions and handling tasks related to office operations.
The third explanation is that occupations within certain sectors often require a high level of skill and involve extensive interpersonal interactions. For instance, in managerial positions (e.g., CEOs), tasks typically encompass responsibilities such as delegating tasks, attending events and meetings, and negotiating contracts - all of which heavily rely on human interpersonal communication. Similarly, in education, a teacher’s role primarily revolves around delivering educational materials in person, a task that demands both physical presence and a higher level of education and specialized knowledge. In the legal sector, impacted occupations include those involved in drafting legal documents or transcribing pretrial and trial proceedings, such as court reporters, in alignment with previous qualitative analyses on legal occupations [31]. Conversely, roles requiring the design of bespoke legal solutions remain unaffected. Notably, client-facing tasks in the legal sector are also not impacted, resulting in an overall expectation of limited impact on the legal sector. Finally, in the financial sector, there are still occupations that necessitate human interactions, such as clerks, sales agents, and tellers.
2.3 Most- and Least-Impacted Geographical Regions
Figure 4: AI Impact on U.S. counties. Orange colors represent higher impact of AI, while purple colors represent lower impact.
Just as with the occupations and industry sectors analyses, the AII measure can be used to study the impact of AI on regional-level outcomes. In fact, patents have been used to study the factors that influence the spread of technology among regional areas [19], as AI is permeating in urban design and planning [32]. Previous research has consistently shown that technology spillovers tend to concentrate within specific geographic regions. This concentration is often attributed to the effective transfer and dissemination of knowledge, a process facilitated by localized interactions involving communication, collaboration, social interactions, and the presence of a local pool of human capital [33, 34].
By calculating the AII measure at the regional level, it became evident that certain states, and even at a more detailed county-level granularity (Figure 4), experienced lower levels of impact compared to others. The East Coast, on average, experiences a greater impact of AI compared to the West Coast. However, Washington and California are exceptions to this pattern, as they are highly impacted due to the presence of Seattle and the Bay Area.
However, the extent to which local interactions can yield significant technological advancements depends on regional capabilities that govern innovation processes. Income inequality emerges as a factor of concern, as it has been found to have adverse effects on regional growth with total wages growing superlinearly as the cities increase in size [35]. Wilkinson and Pickett [36] argue that more equal societies, where income and wealth are distributed more fairly among the population, tend to have better outcomes for their citizens. We correlated the income inequality among U.S. households, calculated as the U.S. state’s Gini coefficient (Figure A.2b), with the AII regional-level measure (Figure 5a and Figure A.2a). We found a moderate positive correlation (Pearson’s $r$ = 0.30), indicating that as the impact of AI increases, so does income inequality at the state level. Contrary to this trend, New York, Alaska, Utah, and New Hampshire stand out as outliers. On the one hand, New York has high inequality [37]. On the other hand, Alaska, Utah, and New Hampshire have lower inequality, partially attributed to their higher rates of homeownership. While large urban centers grapple with widening disparities, states such as Utah and Alaska exhibit more equitable income distribution potentially tied to specific social and industry characteristics [38].
Figure 5: Three socio-economic indicators as a function of region-level ($z$-scored) AII (orange dots are more than two standard deviations away): (a) income inequality (states with higher income inequality will be the most impacted); (b) attractiveness to the creative class (states with a larger creative class will be the most impacted); and (c) economic diversification (counties with the highest diversification will be the least impacted).
Another plausible explanation is that the least impacted states may not predominantly rely on the knowledge economy. Richard Florida’s seminal work established a link between regional economic prosperity and the presence of the “creative class”. Florida defined the creative class as individuals engaged in creative and knowledge-based industries such as artists, designers, scientists, engineers, researchers, and professionals in fields reliant on creativity and intellectual capital [39, 40]. Building upon Florida’s conceptualization of the creative class, a creative index serves as a metric for assessing how well cities diversify their workforce in terms of talent, technology, and creativity [41]. His theory posits that cities and regions with a higher concentration of the creative class are more likely to foster innovation and economic growth. Hoyman and Faricy [42] further supported this notion by demonstrating that states with a higher percentage of the population aged twenty-five and over holding a bachelor’s degree or higher in metropolitan statistical areas tend to be more economically successful. Additionally, investments in talent and technology have been shown to predict the retention of the creative class, ultimately contributing to state income growth and equality [43]. Using Florida’s creativity index as a proxy for the knowledge economy, we found a moderate positive correlation (Pearson’s $r$ = 0.40) between the AII regional-level measure and states’ creativity index (Figure A.2c). This implies that states whose economies are primarily built around the creative class, attracting individuals who excel in information technology and healthcare, are more likely to be impacted by AI. Although the presence of the creative class has previously been associated with economic prosperity, regions that overly concentrate on knowledge-based economies are likely to face significant challenges in upskilling and reskilling their workforces. The state of Arkansas, however, is the sole data point exceeding two standard deviations. Arkansas has been making strides in attracting the creative class, an essential component of a modern and dynamic economy Reports indicate that between 2015 and 2017, the state’s arts and cultural production grew by 3% annually, outpacing the growth of the gross state product and showing a significant value added to the economy by the arts and cultural sector [44]. A key factor behind this is the formation of the Arkansas Legislative Arts and Technology Bootcamp through Act 577 [45]. This initiative is focused on expanding the creative economy and aims to increase revenue and attract creative talents to the state by involving legislative leaders in the efforts.
Table 3: Most- and least-impacted counties. The county-level AII is calculated as the employment-weighted average of sector-level AII scores, weighted by the importance of each economic sector in the county. For each county, we also report the economic diversity (higher values indicate greater diversity) and the predominant employment sector. The economic diversity values for all counties shown are less than 2.8, which is the cutoff for the 1st quartile. This indicates that these counties have relatively low economic diversity, with their workforce concentrated in just a few sectors.
Rank Most-impacted Economic Diversity Most Prevalent Sector Least-impacted Economic Diversity Most Prevalent Sector
1 Issaquena County, MS 0.54 Healthcare Wallace County, KS 1.55 Finance and insurance
2 Petroleum County, MT 1.30 Healthcare Hinsdale County, CO 1.55 Construction
3 Buffalo County, SD 1.46 Healthcare Twiggs County, GA 1.00 Education
4 Macon County, AL 1.67 Healthcare Heard County, GA 1.18 Education
5 Jefferson County, MS 1.79 Healthcare Graham County, NC 2.39 Education
6 Starr County, TX 1.74 Healthcare Garfield County, MT 1.24 Education
7 Iron County, MO 1.80 Healthcare Storey County, NV 1.00 Transportation
8 Metcalfe County, KY 1.99 Healthcare Phillips County, MT 2.64 Education
9 Calhoun County, AR 0.55 Manufacturing Clark County, ID 1.07 Education
10 Ripley County, MO 2.27 Healthcare Ottawa County, KS 2.51 Finance and insurance
11 Galax City, VA 1.94 Healthcare Zavala County, TX 1.39 Education
12 Dillingham Census Area, AK 0.90 Manufacturing McPherson County, SD 1.29 Education
13 Evangeline Parish, Louisiana 2.35 Healthcare Red Lake County, MN 2.46 Education
14 Aleutians East Borough, AK 0.69 Manufacturing Judith Basin County, MT 1.99 Education
15 Franklin City, VA 2.37 Healthcare Gates County, NC 1.99 Education
16 Sherman County, NE 2.20 Healthcare Calhoun County, IL 2.11 Education
17 Reynolds County, MO 2.33 Manufacturing Kingman County, KS 2.85 Education
18 Borden County, TX 1.87 Healthcare Clay County, TX 1.65 Education
19 Todd County, SD 1.57 Healthcare Denali Borough, AK 0.91 Accommodation and food services
20 Wayne County, KY 2.14 Manufacturing Lincoln County, KS 2.25 Education
The third explanation lies in the fact that the least impacted states have different regions that are more economically resilient. Previous research has found that increased economic diversity at the regional level promotes greater economic resilience. Davies [46] analyzed regional economies across Europe and found that regions with more diversity in their workforce composition exhibited higher regional stability and faster recovery from the 2008-2010 global financial crisis. Similarly, Dissart [47] found that U.S. regions with greater diversity across sectors saw lower volatility in employment and income growth, indicating higher regional economic resilience. The proposed mechanism in these studies is that diverse regional economies allow for greater adaptability and adjustment to sector-specific economic shocks. When a recession damages an individual sector, a region with diverse industries can absorb the shock better as workers can shift to unaffected sectors more easily [48]. Similarly, one can hypothesize that when a sector is affected due to automation, a region that diversifies its workforce will be more economically resilient. In contrast, specialized regional economies centered around one or a few dominant industries have been found to be more vulnerable to sector-specific disruptions. Martin [49] found that during recessions, regions in Europe with less complex, less diverse economies suffered greater rises in unemployment rates due to their concentration in a small number of industries that were severely impacted (e.g., manufacturing). To investigate whether finer-grained regional differences are present, we zoomed in on our analysis at the county level. This closer examination allowed us to delve into how localized regional capabilities correlate with the impact of AI. In so doing, we correlated a measure of economic diversity (computed as the entropy of workforce distribution within a county, as detailed in §4.4 and shown in Figure A.2d) with the AII county-level measure (Figure 5b). Two distinct groups of counties of resilient economies emerged (Table 3). The first group comprises counties with diversified economies where the workforce is engaged in a wide range of industry sectors. In fact, diversified metropolitan areas tend to experience more economic growth [18]. The second group consists of counties where the workforce specializes in industry sectors that are least impacted such as education. By contrast, counties specialized in most impacted sectors such as manufacturing and healthcare suffer from such a specialization.
3 Discussion
3.1 Consensus and Discrepancies in Previous Literature
Unlike previous methods for assessing the impact of AI on occupations, which either rely on a finite set of abilities (e.g., manual dexterity) linked to specific occupations [4, 14] or employ subjective evaluations to determine tasks’ suitability for automation [3], our AII measure provides an objective approach by leveraging patent data to capture the dynamic landscape of technological advancements in AI.
To position our findings, we explore the consensus or lack thereof of which occupations will be impacted by AI in the literature.
Consistent results in the literature. Frey and Osborne’s [14], AIOE [4], SML [3], and Webb’s [2] concur that low-skilled occupations typically involving human labor are not impacted by AI. These occupations are typically found in industry sectors such as construction, where tasks primarily involve manual labor and the operation of machinery. Our AII measure aligns with these observations by categorizing such occupations as among the least impacted by AI.
Inconsistent results in the literature. While there has been a clear consensus in previous literature regarding manual labor occupations, we have identified four types of occupations for which consensus has been so far lacking. These categories include: occupations requiring basic and non-specialized skills; occupations requiring high skills and interpersonal interaction-based occupations; those where tasks are replaced by AI-enhanced hardware; and those involving extensive data analysis and processing.
Occupations that require basic and non-specialized skills (least impacted according to our approach). Frey and Osborne’s findings [14] indicated that occupations within the accommodation and food services sector (e.g., cooks, dishwashers, waiters, bartenders) are highly-impacted. Similarly, Webb’s findings suggest that non-routine manual occupations are highly impacted; there are patents matching those occupations, which may well be coarse-grained matches, as exemplified in Table C.1.
However, SML [3] and AIOE [4] found the very same occupations to be among the least-impacted. Our approach aligns with SML and AIOE, identifying these occupations as among the least-impacted due to the absence of AI patents automating the manual tasks associated with them.
Occupations that are highly-skilled and involve interpersonal interactions (least impacted according to our approach). AIOE [4] found that occupations requiring high skills and interpersonal interactions, such as those in the education sector (e.g., teachers) or managerial positions (e.g., CEOs), are highly-impacted. Similarly, Webb found that occupations that require interpersonal tasks are hard to automate [2]. In contrast, SML [3] found these occupations, particularly managerial positions, to be among the least impacted, aligning with Frey and Osborne’s method [14]. However, in the case of SML, there was a distinction within occupations involving interpersonal interactions. While SML found managerial positions to be among the least impacted, it identified occupations such as concierges, which also involve interpersonal interactions, as more likely to be impacted by AI. Upon closer examination of the SML method, which relies on rubrics (i.e., a type of scoring guide for crowdworkers to assess the suitability of tasks for machine learning, as shown in Table A.2, we noted that the corresponding definition in the rubrics entailed a “wide range of interactions,” making it challenging to capture their nuances. Additionally, since SML relies on crowdsourced data, there is a potential for subjective bias or a lack of full understanding of the nuances of interpersonal interactions. Finally, interactions were captured by only one item out of the 23 items in the rubrics making that item’s contribution to the overall score limited. Our approach aligns with Frey and Osborne’s [14], Webb’s [2] and, to some extent with SML [3] as no AI patents were found to target these occupations.
Occupations that consist of tasks that are replaced by AI-enhanced hardware (most impacted according to our approach). Frey and Osborne [14] discovered that occupations in healthcare, such as MRI and cardiovascular technologists, which are likely to be replaced by AI-enhanced hardware, were among the least-impacted. In contrast, SML [3] and AIOE [4] identified these same healthcare occupations as among the most-impacted ones. Webb [2] found that occupations involving non-routine manual tasks (e.g., operating devices or equipment in healthcare or manufacturing) to be among the most-impacted ones. Our approach aligns with SML, AIOE, and Webb’s method due to the presence of AI patents automating tasks using AI-enhanced hardware. For example, in healthcare, patents have automated tasks such as medical imaging and diagnosis using machine learning models, the development of treatment plans, the creation of medical devices, the recording and analysis of patient data, and even the prediction of optimal radiation therapy doses, dental treatment plans, and the processing of medical data.
Occupations that require extensive data analysis and processing (most impacted according to our approach). Frey and Osborne [14] found that occupations that require extensive data analysis and processing, such as those in information technology or manufacturing (e.g., software developers, chemists, aerospace engineers), were classified as among the least-impacted. In contrast, SML [3] and AIOE [4] categorized these very same occupations as among the most-impacted. Similarly, Webb [2] identified that occupations involving non-routine cognitive analytic tasks to be among the most-impacted ones. Our findings align with SML, AIOE, and Webb’s method due to the presence of patents related to tasks that require data analysis and processing. For example, patents about training robots for task execution, image and video processing, speech recognition, and natural language processing, as well as the integration of machine learning into software systems. These patents automate tasks such as troubleshooting networks, code reviews, optimizing supply chain logistics, planning material dumping operations, and supporting maintenance, including determining aircraft conditions.
3.2 Implications
The impact of AI on occupations carries important implications for the workforce. However, when placed our findings within the context of previous literature, it became evident that there is a lack of consensus regarding which occupations will be affected and which will remain unaffected. As an initial step, we contend that achieving consensus is crucial to start formulating effective policies to address the ongoing transformations in the labor market. To initiate this process, with our findings in mind, we outline three key areas in which initiatives can be developed.
Initiatives for White-Collar Workers. Policymakers and employers should launch specific initiatives targeting white-collar occupations in sectors such as information technology, manufacturing, and healthcare. These initiatives can equip workers with the skills needed for high-value, creative, and complex tasks. For example, a manufacturing worker can undergo training in robotics programming, enabling them to effectively operate and maintain AI-driven machinery. Similarly, healthcare professionals can acquire telemedicine and data analytics skills to enhance patient care and diagnostics.
Initiatives for Blue-Collar Workers. Blue-collar occupations, predominantly found in agriculture, accommodation and food services, and construction, typically involve low-skilled work demanding physical labor. While previous literature has suggested reallocating low-skilled workers to tasks less susceptible to computerization [14], such as those requiring creative and social intelligence, we argue that (re)skilling and upskilling [10] should be approached cautiously. Our regional analysis emphasizes the importance of striking a balance in upskilling low-skilled workers for roles demanding creative and social intelligence. This is particularly important because regions that excessively rely on knowledge-based economies are likely to face significant AI impact.
Initiatives for Continuous Learning and Interdisciplinary Training. Promoting a culture of continuous learning and skill development is essential. Employers can encourage individuals to embrace lifelong learning through online courses, certifications, and vocational training, enabling them to adapt to the changing occupational landscape [50]. Similarly, encouraging interdisciplinary training can prepare the workforce for the demands of AI-augmented occupations. For example, blending traditional engineering with AI and machine learning training can create a workforce capable of developing and maintaining AI-enhanced systems across sectors.
3.3 Limitations and Future Work
This work has six primary limitations. First, our measure reflects whether a patent impacts a task but cannot distinguish whether the impact means replacing the task or augmenting it, which has been a long standing debate in the literature [23, 51, 4, 52, 53].
Second, our analysis is based on currently existing tasks, capturing their potential automation, and cannot be used to study the emergence of new tasks as a result of that automation (referred to in previous work as ‘work augmentation’ [23]).
Third, our method relies on patent abstracts to provide a finer-grained understanding of occupational tasks, mainly from the US-focused USPTO dataset. Future research should explore potential cultural differences in how patents are written and used in other contexts, such as the European Union.
Fourth, it is important to acknowledge that the existing portfolio of USPTO patents may not comprehensively cover all the innovations that may impact a particular occupation. To address this limitation, one potential approach could involve supplementing patent data with other sources, including research papers and code repositories [54].
Fifth, our assumption is that the likelihood of an AI system being built is determined by whether it is patented. While generally true, there are exceptions. A patented system may not be built, as the patent could be intended for defensive or offensive purposes [55, 56]. Conversely, a non-patented system might still be constructed, with its design protected by secrecy or trademarks [57, 58]. Furthermore, patenting rates vary among sectors [59], and there is a lag between an innovation being patented and its use and impacts diffusing across the economy [54].
Finally, our analysis is based solely on a concise yet comprehensive five-year time window, spanning from 2015 to 2020, with patents from these years accounting for 92% of the total AI patents granted until 2020. We also repeated the analysis for 2010-2020 and found no significant difference in the results (as shown in Table D.1 and Figure D.1). This does not capture emerging technologies, such as Large Language Models (LLMs). Future research could replicate our methodology to assess the impact of emerging technologies, such as cryptocurrency, the metaverse, and LLMs by using upcoming patents in those fields.
4 Materials and Methods
Figure 6: Framework automating AI Impact (AII) measure. Using the Sentence-T5 (ST5) model, we first generated two vector representations (embeddings): one capturing the semantic meaning of an AI patent, and the other the semantic meaning of a task description. We then computed the patent-task cosine similarity from the embeddings on all patent-task pairs. This process was conducted to identify which tasks were impacted by which AI patents. Finally, for each occupation, we calculated the proportion of impacted tasks out of the total ones, and this proportion determined the AII of that occupation.
4.1 Datasets
Occupation Dataset. We collected detailed task descriptions for a wide range of occupations from the O*NET database [60], a widely used source in occupational studies [61, 62, 3, 4, 63]. In total, we collected 773 unique occupations and 19,498 unique tasks from O*NET 24.3 version released in May 2020. The distribution of these tasks ranges from 4 to 286, with a median of 20 tasks per occupation (Figure A.1).
Patent Dataset. To obtain a corpus of AI patents, we first retrieved 74,875 patents granted by the United States Patent and Trademark Office (USPTO) between 2015-2020 that were classified to be about AI based on an official way of coding patents called PATENTSCOPE AI Index [64] to then filter away patents only tangentially related to AI. We selected the subset of patents in the index class core AI applications (Table A.1). This resulted in a final corpus of 12,984 AI patents.
4.2 Measuring AI Impact (AII) on Occupation Tasks
Sentence-Transformers. We developed a Sentence-Transformers Deep Learning framework [65] for Natural Language Processing that uses the Sentence-T5 (ST5) architecture [26] to convert input text into “semantic vector representations” called “embeddings”. These embeddings capture the semantic information of the text and allow us to mathematically compute the similarity of a pair of text snippets.
AI Impact (AII). We defined AII as a measure of “the exposure to AI by measuring the extent to which an occupation’s tasks are associated with patents”. For each task, we identified the patent with the highest task-patent similarity score (Equation 1) to represent the AI impact $\alpha_{t}$ on task $t$ indicating the extent to which task $t$ aligns with AI-related innovations:
$\alpha_{t}=\max_{p}\text{sim}(v_{t},v_{p})$ (1)
where $\text{sim}(v_{t},v_{p})$ is the cosine similarity between the embeddings of task $t$ and the embeddings of patent $p$. We computed the impact of AI on task $t$ by taking the maximum similarity value. We took the maximum instead of, say, the average because if the average was used, similarity scores from patents that are not particularly relevant to the task would be factored into the calculation, thereby diluting the AI impact score (Figure B.1).
4.3 Measures of AII on Occupations, Industry Sectors, and Regions
AI Impact on Occupations. We computed the AI impact $y_{j}$ on occupation $j$ by computing the number of AI-impacted tasks over the total number of $j$’s tasks:
$y_{j}=\frac{\sum_{t\in\text{tasks}(j)}1_{\alpha_{t}>p_{90}(\alpha)}}{\sum_{t\in\text{tasks}(j)}1}$ (2)
where $p_{90}(\alpha)$ is the 90th percentile of AI impact values computed on all occupations’ tasks, and $1_{\alpha_{t}>p_{90}(\alpha)}$ is an indicator function whose value is 1, if $\alpha_{t}>p_{90}(\alpha)$, and 0 otherwise. Using the 90th percentile as the threshold makes the AI impact measure more robust to noise, which was also suggested in a previous study [21]. Given that every task is assigned a similarity value in the previous step, the patent deemed most similar for a specific task might still be unrelated to that task. On the other hand, a higher 95th percentile threshold would be too strict, as 55% of the occupations would have zero impacted tasks. To further validate our task-patent matching method, two authors independently assessed the relevance of a patent to a task in a random sample of 100 task-patent pairs. Overall, their agreement was nearly perfect, with a Cohen’s Kappa of 0.84.
AI Impact on Industry Sectors. To determine the impact of AI on industry sector $s$, we calculated the mean AII score across all occupations $j$ associated with sector $s$ (Equation 3). Occupation $j$ was assigned to sector $s$, if over 50% of workers in $j$ were employed in $s$:
$\pi_{s}=\frac{1}{N_{s}}\sum_{j\in\text{occupations}(s)}y_{j}$ (3)
where $N_{s}$ is the number of occupations associated with $s$. If more than half of the workers in an occupation are employed in a particular sector, it can be reasonably concluded that this occupation is primarily associated with that sector [4]. Lower thresholds might lead to occupations being associated with multiple sectors, making the results less specific. Conversely, a higher threshold might be too restrictive, potentially excluding occupations with a significant presence in a sector, even if not overwhelmingly so.
AI Impact on Regions. To determine regional AII scores, we used the National Occupational Employment and Wage Estimates dataset published in May 2022 by the U.S. Bureau of Labor Statistics. For each region, we calculated a weighted average of the AII scores across sectors, weighting each sector by the number of employees in that sector within the region.
$\Omega_{r}=\frac{\sum_{s\in\text{sectors}(r)}\pi_{s}w_{rs}}{\sum_{s\in\text{sectors}(r)}w_{rs}}$ (4)
where $w_{rs}$ is the number of employees associated with sector $s$ within region $r$. $\Omega_{r}$ is $r$’s regional AII score that accounts for both the AI impact on occupations in that region and the distribution of the workforce across sectors. Sectors that employ more people in a region have a larger influence on that region’s overall labor market dynamics. Weighting by the number of employees ensures that sectors with a higher workforce representation influence the regional AII score proportionately. Moreover, some sectors might have extreme AII scores (either very high or very low). A simple average could be skewed by these outliers. By weighting sectors based on the number of employees, the measure ensures that the regional AII score is not disproportionately affected by these smaller sectors. This method of weighted averaging aligns with methodologies in previous studies [4, 3, 2].
4.4 Measuring Economic Diversity
To measure the economic diversity of a region, we collected its workforce distribution data from the National Occupational Employment and Wage Estimates dataset published in May 2022 by the U.S. Bureau of Labor Statistics. For each region, the economic diversity is computed as the entropy of the workforce distribution:
$H(r)=-\sum_{s\in\text{sectors}(r)}\frac{w_{rs}}{N_{r}}\log\frac{w_{rs}}{N_{r}}$ (5)
where $w_{rs}$ is the number of employees associated with sector $s$ within region $r$ and $N_{r}$ is the total number of employees within region $r$.
4.5 Thematic Analysis on Occupations and Industry Sectors
Occupations. To identify emergent themes that characterize least or most impacted occupations, we conducted a thematic analysis [66, 67] on the task-patent pairs for all tasks associated with those occupations. This process consists of two steps: open coding, in which textual data is broken up into discrete parts and then coded; and axial coding, in which the researcher draws connections between the generated codes [68]. We first applied open coding to identify key concepts that emerged across multiple task-patent pairs; specifically two of the authors read all task-patent pairs, and marked them with keywords that reflected the key concepts expressed in the text. They then used axial coding to identify relationships between the most frequent keywords to summarize them in semantically cohesive themes (e.g., healthcare, information technology, and manufacturing). Themes were reviewed in a recursive manner rather than linear, by re-evaluating and adjusting them as new task-patent pairs were parsed.
Industry Sectors. Just as with occupations, we conducted thematic analysis [66, 67, 68] to uncover potential explanations for the most- and least-impacted sectors. We again applied open coding to identify key concepts that emerged on the descriptions of occupations (examples presented in Table A.3) across the industry sectors under study, with the same two authors reading all descriptions and marking them with keywords that reflected the key concepts expressed in the text. We then again used axial coding to identify relationships and potential explanations.
Supplementary Information
Appendix A Data
A.1 Data Collection
In a total of 12,984 AI patents granted between 2015 and 2020, the majority contained the keyword machine learning (46%), followed by the keyword neural network (32%), artificial intelligence (9%), and deep learning (6%) (Table A.1).
Table A.1: Number of patents based on keywords.
Keywords Number of patents
machine learning 6084
[…]
From 774 occupations defined in O*NET, the number of tasks for each occupation ranges from 4 to 286, with a median of 20 tasks (Figure A.1).
Figure A.1: The distribution of the number of tasks per occupation
To compute SML scores, Brynjolfsson et al. [3] used a set of rubrics to guide crowdworkers to assess the suitability of tasks for machine learning (Table A.2). Note that only one item in the rubrics is about interpersonal interactions.
Table A.2: Rubrics to compute SML scores as developed by Brynjolfsson et al. [3].
No
Rubric
Q1
Information needed to complete the task (inputs) and outputs can be explicitly specified in machine-readable format
[…]
Task information is recorded or recordable by computer
[…]
Task does not require a wide range of complex outputs (mental and/or physical)
[…]
Task feedback (on the success of outputs) is immediate
[…]
The task output is error tolerant
[…]
It is not important that outputs are perceived to come from a human
[…]
Task does not require complex, abstract reasoning
[…]
Task is principally concerned with matching information to concepts, labels, predictions, or actions
[…]
Task does not require detailed, wide-ranging conversational interaction with a customer or other person
[…]
Task is highly routine and repeated frequently
[…]
Task is describable with rules
[…]
There is no need to explain decisions during task execution
[…]
Task can be converted to answering multiple choice questions, ranking alternative options, predicting a number, or grouping similar objects
[…]
Long term planning is not required to successfully complete the task
[…]
The task requires working with text data or might require working with text in the future
[…]
The task requires working with image/video data or might require working with image/video data in the future
[…]
The task requires working with speech data or might require working with speech data in the future
[…]
The task requires working with other types of data (other than text, image/video, and speech)
[…]
Many components of the task can be completed in a second or less
[…]
Each instance, completion, or execution of the task is similar to the other instances in how it is done and these actions can be measured
[…]
Actions in the task must be completed in a very specific order, and practicing the task to get better is easy
To conduct the thematic analysis on industry sectors, we used the occupation titles and descriptions in O*NET. For brevity, we provide 10 examples in Table A.3 and we refer the reader to the full list in O*NET [60].
Table A.3: Examples of occupational descriptions in O*NET.
Occupation Title
Sector
Description
Cardiovascular Technologists and Technicians
Healthcare
Conduct tests on pulmonary or cardiovascular systems of patients for diagnostic purposes. May conduct or assist in electrocardiograms, cardiac catheterizations, pulmonary functions, lung capacity, and similar tests. Includes vascular technologists.
[…]
Healthcare
Examine, diagnose, and treat dental malocclusions and oral cavity anomalies. Design and fabricate appliances to realign teeth and jaws to produce and maintain normal function and to improve appearance.
[…]
Healthcare
Compile, process, and maintain medical records of hospital and clinic patients in a manner consistent with medical, administrative, ethical, legal, and regulatory requirements of the health care system. Process, maintain, compile, and report patient information for health requirements and standards in a manner consistent with the healthcare industry’s numerical coding system.
[…]
Manufacturing
Develop programs to control machining or processing of metal or plastic parts by automatic machine tools, equipment, or systems.
[…]
Information
Create special effects, animation, or other visual images using film, video, computers, or other electronic tools and media for use in products or creations, such as computer games, movies, music videos, and commercials.
[…]
Healthcare
Operate Magnetic Resonance Imaging (MRI) scanners. Monitor patient safety and comfort, and view images of area being scanned to ensure quality of pictures. May administer gadolinium contrast dosage intravenously. May interview patient, explain MRI procedures, and position patient on examining table. May enter into the computer data such as patient history, anatomical area to be scanned, orientation specified, and position of entry.
[…]
Healthcare
Prepare, administer, and measure radioactive isotopes in therapeutic, diagnostic, and tracer studies using a variety of radioisotope equipment. Prepare stock solutions of radioactive materials and calculate doses to be administered by radiologists. Subject patients to radiation. Execute blood volume, red cell survival, and fat absorption studies following standard laboratory techniques.
[…]
Information
Develop, create, and modify general computer applications software or specialized utility programs. Analyze user needs and develop software solutions. Design software or customize software for client use with the aim of optimizing operational efficiency. May analyze and design databases within an application area, working individually or coordinating database development as part of a team. May supervise computer programmers.
[…]
Manufacturing
Operate, test, maintain, or calibrate unmanned, automated, servo-mechanical, or electromechanical equipment. May operate unmanned submarines, aircraft, or other equipment at worksites, such as oil rigs, deep ocean exploration, or hazardous waste removal. May assist engineers in testing and designing robotics equipment.
[…]
Manufacturing
Operate industrial trucks or tractors equipped to move materials around a warehouse, storage yard, factory, construction site, or similar location.
A.2 Summary Statistics
We checked the four variables’ distribution to ensure that they were normally distributed before using them in our regional-level analysis. All but economic diversity were normally distributed. Since we did not conduct any correlation analysis on the economic diversity, a log transformation was not needed. Instead, we plotted the quadratic fit of this variable to understand its relationship with the AII measure computed at the regional level.
Figure A.2: Frequency distribution plots of: (a) the AII measure; (b) income inequality; (c) creativity index; and (d) economic diversity.
Appendix B Task-Patent Matching
B.1 Validating the similarity metric used for matching
To ensure the specificity of our task-patent matching method, we plotted the mean and max of the AII measure. We empirically found that by using the max, we retained the most relevant patent (Figure B.1). To further validate our task-patent matching method, we manually annotated a random sample of 100 task-patent pairs in terms of task-patent relevance; two of the authors who did the annotations achieved a Cohen’s kappa of 0.84 (Kappa values in the [0.81, 1] range suggest almost perfect agreement).
Figure B.1: Distributions of the task-patent similarity based on different aggregations. By focusing only on the most relevant patents, we captured the specificity of the AI impact measurement for each task, reducing noise, and incorporating a wider range of AI-related advancements.
B.2 Matching abstracts (not only titles)
Patents tend to be general, and specific tasks may not be mentioned simply in the patent title. To address this, our method uses the patent abstract that is richer in text and more likely to reference a specific task or even entire sector(s) (Table B.1). As a result, we were able to match the same patent with occupations in multiple sectors. For example, a patent that provides methods for radiotherapy treatment planning not only matches the task of radiation therapists, but also counselors (abuse and behavioral disorder) who need treatment plans. Another example is a patent for custom-fitting a service solution to consumer requirements that is applicable to occupations from different sectors (e.g., software developers, social workers, and solar photovoltaic installers).
Table B.1: Example task-patent pairs matched with AII. The dash (-) sign means that the patent is the same as the cell above it.
Occupation
Sector
Task
Matching Patent Title
Matching Patent Abstract
Similarity
Substance Abuse and Behavioral Disorder Counselors
Healthcare
Develop client treatment plans based on research, clinical experience, and client histories.
Methods and systems for radiotherapy treatment planning
Example methods for radiotherapy treatment planning are provided. One example method may include obtaining training data that includes multiple treatment plans associated with respective multiple past patients; […]
0.8018
[…]
Healthcare
Develop and implement treatment plans based on clinical experience and knowledge.
0.8082
[…]
Healthcare
Administer prescribed doses of radiation to specific body parts, using radiation therapy equipment according to established practices and standards.
0.8330
Architectural and Civil Drafters
Manufacturing
Reproduce drawings on copy machines or trace copies of plans and drawings, using transparent paper or cloth, ink, pencil, and standard drafting instruments.
Line drawing generation
Computing systems and computer-implemented methods can be used for automatically generating a digital line drawing of the contents of a photograph. […] The training data set teaches the neural network to trace the edges and features of objects in the photographs, as well as which edges or features can be ignored. […]
0.8038
[…]
Information
Create pen-and-paper images to be scanned, edited, colored, textured, or animated by computer.
0.8396
[…]
Public administration
Create sketches and diagrams by hand or with computer software to depict crime scenes.
0.8195
[…]
Other services
Operate scanners or related computer equipment to digitize negatives, photographic prints, or other images.
0.8167
Software Developers, Applications
Information
Consult with customers about software system design and maintenance.
System and method for custom-fitting services to consumer requirements
Systems and methods for custom-fitting a service solution to consumer requirements are provided. […]
0.8020
[…]
Education
Interview clients individually, in families, or in groups, assessing their situations, capabilities, and problems to determine what services are required to meet their needs.
0.8010
[…]
Construction
Determine photovoltaic (PV) system designs or configurations based on factors such as customer needs, expectations, and site conditions.
0.8031
Appendix C Previous Attempts to Link Tasks to Patents
Word-matching methods linking tasks to patents. A word-matching method employs a dictionary approach, parsing task descriptions to identify verb-noun pairs associated with each task [2]. This approach captures the task essence concisely and specifically, such as the pair “install, sensor”, offering a clear representation of the task. Using the same method, AI patent titles are also processed to extract verb-noun pairs describing the tasks targeted by each patent. The relative frequency of similar pairs in tasks and patent titles determines the AI exposure score, indicating the level of task exposure to AI.
Word-matching methods: Potential false positives. Webb method [2] identified “elevator installers” as one of the most impacted occupations by AI. However, he also acknowledged that this classification is an example of false positives caused by word matching methods being coarse-grained. By applying our deep-learning method instead, we learn that nearly all the tasks of the occupation “elevator installers” are indeed not impacted (Table C.1): only 1 out of 20 tasks was impacted. This speaks to the robustness of a deep-learning approaches compared to a word-matching approach.
Table C.1: List of the the tasks for the “elevator installers and repairers” occupation, and whether each task is impacted or not based on AII. A task is impacted when there is a patent with a similarity exceeding the 90th percentile threshold.
Task Is impacted?
Locate malfunctions in brakes, motors, switches, and signal and control systems, using test equipment.
Yes
[…]
No
[…]
No
[…]
No
[…]
No
[…]
No
[…]
No
[…]
No
[…]
No
[…]
No
[…]
No
[…]
No
[…]
No
[…]
No
[…]
No
[…]
No
[…]
No
[…]
No
[…]
No
[…]
No
Word-matching methods: Potential false negatives. In addition to, at times, matching incorrectly tasks and patents (false positives), word matching methods may also fail to return genuine matches (false negatives). For example, Webb’s paper reported a variety of tasks for which no patent was found (Table C.2). However, by applying AII to those tasks, we instead found that there are indeed matching patents. This discrepancy likely arises from word-matching methods discarding multiple verb-noun pairs, and overlooking details in patent abstracts.
Table C.2: List of tasks with zero AI Exposure Score, indicating no associated patents [2]. In contrast, AII identifies relevant patents for those tasks.
Task
Extracted pairs AI exposure score
Most Similar Patent Abstract
Similarity
Document and maintain records of precision agriculture information.
(maintain, record) 0.000
A method and system for predicting soil and/or plant condition in precision agriculture with a classification of measurement data for providing an assignment of a measurement parcel to classes of interest. The assignment is used for providing action recommendations, particularly in real time or close to real time, to a farmer and/or to an agricultural device based on acquired measurement data, particularly remote sensing data, and wherein a classification model is trained by a machine learning algorithm, e.g. relying on deep learning for supervised and/or unsupervised learning, and is potentially continuously refined and adapted thanks to a feedback procedure.
0.790
[…]
(apply, information) 0.000
A method and system for predicting soil and/or plant condition in precision agriculture with a classification of measurement data for providing an assignment of a measurement parcel to classes of interest. The assignment is used for providing action recommendations, particularly in real time or close to real time, to a farmer and/or to an agricultural device based on acquired measurement data, particularly remote sensing data, and wherein a classification model is trained by a machine learning algorithm, e.g. relying on deep learning for supervised and/or unsupervised learning, and is potentially continuously refined and adapted thanks to a feedback procedure.
0.813
[…]
(maintain, sensor) 0.000
A vehicle computing system validates location data received from a Global Navigation Satellite System receiver with other sensor data. In one embodiment, the system calculates velocities with the location data and the other sensor data. The system generates a probabilistic model for velocity with a velocity calculated with location data and variance associated with the location data. The system determines a confidence score by applying the probabilistic model to one or more of the velocities calculated with other sensor data. In another embodiment, the system implements a machine learning model that considers features extracted from the sensor data. The system generates a feature vector for the location data and determines a confidence score for the location data by applying the machine learning model to the feature vector. Based on the confidence score, the system can validate the location data. The validated location data is useful for navigation and map updates.
0.817
Appendix D Most- and Least-Impacted Occupations and Industry Sections
D.1 Between 2010 and 2020
We computed the AII measure on occupations and industry sectors in a larger time window between 2010 and 2020. Similarly, we observed that the most impacted occupations come primarily from healthcare, information technology, and manufacturing, whereas the least impacted ones come from finance and insurance, education, and construction.
Table D.1: 20 most- and least-impacted occupations ranked by the AII (Artificial Intelligence Impact) measure using AI patents from 2010 to 2020. For the 20 most- and least-impacted occupations, only 1 entry differs for each. This is because patents in 2015-2020 account for 96% of the total patents in 2010-2020.
Rank Most-impacted Least-impacted
1 Orthodontists Pile-Driver Operators
[…]
Figure D.1: Overall AI impact on industry sectors from 2010-2020, ranked from highest to lowest sector-level AII scores. A sector-level AII score represents the mean AII across all occupations within that sector. The only difference from the 2015-2020 version is the ordering of other services and construction sectors.
D.2 Between 2010 and 2022
We computed the AII measure on occupations and industry sectors in a larger time window between 2010 and 2022. In this extended timeframe, we again observed that the most impacted occupations primarily come from healthcare, information technology, and manufacturing. Additionally, we noted an increasing impact of AI in scientific occupations.
Table D.2: 20 most- and least-impacted occupations ranked by the AII (Artificial Intelligence Impact) measure using AI patents from 2010 to 2022.
Rank Most-impacted Least-impacted
1 Cardiovascular Technologists and Technicians Pile-Driver Operators
[…]
Figure D.2: Overall AI impact on industry sectors from 2010-2022, ranked from highest to lowest sector-level AII scores. A sector-level AII score represents the mean AII across all occupations within that sector.